there are 3 pumps that fill a pool the first pump can fill the pool in 3 hours the second pump can fill the pull in 4 hours and the 3rd pump can fill the pool in 6 hours how long would it take to fill the pool if all 3 pumps are running at the same time
Answer by addingup(3677)

To compare apples with apples, first you have to bring them to a common number. We'll make them all 1 hour. The first take 3 hours, so in 1 hour it fills 1/3 of the pool. The second takes 4 hours, so in 1 hour it fills 1/4 of the pool. And the last is 1/6. Working together, add them up:
1/3x+1/4x+1/6x = 1 Multiply all sides by 12 in order to have 1 as a denominator:
4x+3x+2x= 12
9x= 12
x= 12/9 divide both by 3:
x= 4/3 = 1 1/3 It takes the three pumps working together 1 1/3 hours to fill the pool. An hour has 60 minutes, so 1/3 is 20. Thus, it takes the three pumps 1 hour and 20 minutes to fill the pool.